OpenStudy (anonymous):

Find this value if f(x) = 5/x+2. f (m - 2)

OpenStudy (lalaly):

f(m-2) = 5/(m-2)+2 = 5/m

hero (hero):

\[f(m-2) = \frac{5}{m-2+2} = \frac{5}{m}\]

hero (hero):

Just making it a little clearer.